The Hamiltonian for a spin-1/2 particle at rest is given by H=E0(σz+ασx), where σx and σz are Pauli spin matrices and E0 and α are constants. The eigenvalues of this Hamiltonian are
1
Expert's answer
2020-10-16T11:01:36-0400

Solution

Hamiltonian is given by

"H=E_0(\sigma_z+\alpha\sigma_x)"

Using pauli matrices


"H=E_0\begin{pmatrix}
 1 & 0\\
 0 & -1
\end{pmatrix}++ E_0 \alpha\begin{pmatrix}
 0 & 1 \\
 0 & 1
\end{pmatrix}"

"H=E_0\begin{pmatrix}
 1 & \alpha \\
 \alpha & -1
\end{pmatrix}"

if eigenvalue is "\lambda" then

"H-\lambda I=0"

"E_0\begin{pmatrix}
 (1-\lambda) & \alpha\\
 \alpha & -(1+\lambda)
\end{pmatrix}=0"


"\lambda=+E_0\sqrt{1+\alpha^2}"

or "\lambda=-E_0\sqrt{1+\alpha^2}"